OpenStudy (anonymous):

Hey guys if you can help that would be great!! There are Eight rubies, ten emeralds, and a hundred pearls, which are in thy ear-ring, my beloved, were purchased by me for thee at an equal amount; and the sum of the prices of the three sort of gems was three less than half a hundred; tell me the price of each, auspicious woman. What is the price of one pearl ???? , the price of one emerald ????? , and the the price of one ruby?????

OpenStudy (anonymous):

8r=10E=100P 8r+10E+100P=100/2-3 8r+8r+8r=47 24r=47 r=1.9583333... which is now refered to as A 8r=10E 8(A)=10E 1.566666...=E 8A=100P .1566666...=P (sorry couldnt make a new line)

OpenStudy (anonymous):

r=rubies E=emeralds P=pearls